OpenStudy (anonymous):

How to prove P[(A∪B)∩(¯A∪¯B)] = P(A) + P(B) - 2P(A∩B) via probability manipulations?

OpenStudy (blockcolder):

\[P((A\cup B)\cup(A'\cup B'))=P(A\cup B)+P(A' \cup B')-P((A \cup B)\cap(A'\cup B'))\\ P((A\cup B)\cup(A'\cup B'))=P(A \cup A' \cup B\cup B')=P(U)=1\\ P(A\cup B)=P(A)+P(B)-P(A\cap B)\\ P(A'\cup B')=P((A\cap B)')=1-P(A\cap B)\] Put these all together and you get \[1=P(A)+P(B)-P(A\cap B)+1-P(A\cap B)-P((A \cup B)\cap(A'\cup B'))\\ \Rightarrow P((A \cup B)\cap(A'\cup B'))=P(A)+P(B)-2P(A\cap B)\]